# Distributionally Robust Markov Games with Average Reward

### v2

@@ C.2 Hemi-continuity
Consider a policy sequence $\{\pi_{n}\}\rightarrow\pi$, and the sequence of induced transition matrices $\{\mathsf{P}_{\pi_{n}}\}$. It holds that $\mathsf{P}_{\pi_{n}}^{\star}\rightarrow\mathsf{P}_{\pi}^{\star}$ uniformly on $\mathcal{P}$.
@@
See Prop 8.4.6 in (puterman1994markov) for the proof. ∎
@@
Let $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ be a joint policy of all other agents except agent $i$. Define the operator $\mathcal{T}^{\pi}_{\mathcal{P}}:\mathbb{R}^{\mathcal{S}}\rightarrow\mathbb{R}^{\mathcal{S}\mathcal{A}_{i}}$ as $\mathcal{T}^{\pi}_{\mathcal{P}^{a}_{s}}(h)(s,a)=r^{\pi}(s,a)+\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)$, where $(\mathcal{P}^{\pi})^{a}_{s}=\big\{q(\cdot)=\sum_{b\in\mathcal{A}_{-i}}\pi(b)\mathsf{P}(\cdot|s,(a,b))\ :\ \mathsf{P}(\cdot|s,(a,b))\in\mathcal{P}^{(a,b)}_{s}\big\}$ and $r^{\pi}(s,a)=\sum_{b\in\mathcal{A}_{-i}}\pi(b)r(s,a,b).$ Consider the sequences $\{\pi_{n}\}\rightarrow\pi$ and $\{h_{n}\}\rightarrow h$, then $\mathcal{T}^{\pi_{n}}_{\mathcal{P}}(h_{n})\rightarrow\mathcal{T}^{\pi}_{\mathcal{P}}(h)$ uniformly on $\mathcal{S}\times\mathcal{A}_{i}$.
@@
C.3 Technical Lemmas
In this section, we provide some supporting lemmas necessary for our main analysis.
Lemma C.6.
Define the ergodicity coefficient of a stochastic matrix $\mathsf{P}$ as
Proof.
For any two states $i,j$:
$=(1-\alpha(\mathsf{P}))\text{sp}(V)$
Lemma C.7 (Difference of Robust Operators).
For any $V_{1},V_{2}\in\mathbb{R}^{|\mathcal{S}|}$, there exists a valid transition kernel $\mathsf{P}^{\star}$ constructible from $\mathcal{P}$ such that $\sigma^{\pi}(V_{1})-\sigma^{\pi}(V_{2})=\mathsf{P}^{\star}(V_{1}-V_{2})$.
Proof.
@@ Appendix D Weakly Communicating DR-MGs
Robust average reward and best responses. For a stationary policy $\pi_{i}:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ of agent $i$, define
$g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}\ \triangleq\ \min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\pi_{i}}_{P},\qquad\text{BR}_{i}(\pi_{-i})\ \triangleq\ \Bigl\{\pi_{i}:\ g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}=\sup_{\mu:\,\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\mu}_{P}\Bigr\}.$ (61)
Constant-gain robust Bellman equations. Under ˜2, there exists a constant-gain solution $(u^{\star},\alpha^{\star})$ to the statewise equations (Theorem 2 in (wang2025bellman)):
$u^{\star}(s)$ $=\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\ \sup_{\varphi\in\Delta(\mathcal{A}_{i})}\sum_{a_{i}}\varphi(a_{i})\Big(r_{i}^{\pi_{-i}}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Big),$ (62)
$u^{\star}(s)$ $=\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\ \inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\sum_{a_{i}}\varphi(a_{i})\Big(r_{i}^{\pi_{-i}}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Big),\qquad s\in\mathcal{S},$ (63)
with attainment of both $\inf$ and $\sup$.
@@ D.1 Convexity
For $s\in\mathcal{S}$ define
$L_{s}\!\left(\varphi,p_{s}\right)$ $\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\big),\quad\varphi\in\Delta(\mathcal{A}_{i}),\ p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s},$ (64)
$F_{s}(\varphi)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}L_{s}\!\left(\varphi,p_{s}\right),\qquad\Psi_{s}(p_{s})\triangleq\sup_{\varphi\in\Delta(\mathcal{A}_{i})}L_{s}\!\left(\varphi,p_{s}\right).$ (65)
Then $F_{s}$ is concave and upper hemi-continuous on the compact convex set $\Delta(\mathcal{A}_{i})$ (infimum of continuous affine maps). Using (62)–(63),
@@
$\mathcal{M}_{s}\triangleq\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F_{s}(\varphi).$ (67)
Then $\mathcal{M}_{s}$ is nonempty, convex, compact (boyd2004convex).
@@
By Berge’s Maximum Theorem (berge1877topological), $\Psi_{s}$ is continuous on compact $(\mathcal{P}^{\pi_{-i}})_{s}$, so a minimizer $p_{s}^{\star}$ exists. Then (66) gives $\Psi_{s}(p_{s}^{\star})=u^{\star}(s)$. For any $\varphi\in\mathcal{M}_{s}$,
@@
For $\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ define $\delta_{\nu}(s)\triangleq u^{\star}(s)-F_{s}(\nu(\cdot\mid s))\geq 0$. Under $(\nu,p^{\star})$,
$\limsup_{n\to\infty}\mathbb{E}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}\big(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\big)\right]$ $\leq 0,$ (70)
$\liminf_{n\to\infty}\mathbb{E}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}\big(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\big)\right]$ $\geq-\,\limsup_{n\to\infty}\mathbb{E}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\right].$ (71)
If $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for all $s\in C_{p^{\star}}$, then $g^{\nu}_{p^{\star}}=\alpha^{\star}$.
@@
From Lemma˜D.1, $u^{\star}(s)=\sup_{\varphi}L_{s}\!\left(\varphi,p_{s}^{\star}\right)$, hence
$u^{\star}(S_{t})\ \geq\ \mathbb{E}\!\big[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\big],$ (72)
and telescoping yields the first inequality. Also $F_{s}(\nu(\cdot\mid s))\leq L_{s}\!\left(\nu(\cdot\mid s),p_{s}^{\star}\right)$ implies
$u^{\star}(S_{t})-\delta_{\nu}(S_{t})\ \leq\ \mathbb{E}\!\big[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\big],$ (73)
and telescoping yields the second. If $\delta_{\nu}=0$ on $C_{p^{\star}}$ then its time-average vanishes because $\mathcal{S}\setminus C_{p^{\star}}$ is transient under $(\nu,p^{\star})$, so $g^{\nu}_{p^{\star}}=\alpha^{\star}$. ∎
Proposition D.1.
Let $\text{BR}_{i}(\pi_{-i})$ be the set of stationary best responses (optimal stationary policies of the induced DR-MDP). Then:
  • (a)
    (Necessary) If $\nu\in\text{BR}_{i}(\pi_{-i})$, then $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for every $s\in C_{p^{\star}}$.
  • (b)
    (Sufficient) If $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for all $s\in\mathcal{S}$ (i.e., $\nu$ is statewise greedy for $(u^{\star},\alpha^{\star})$), then $\nu\in\text{BR}_{i}(\pi_{-i})$.
@@
(a) If $\nu(\cdot\mid s)\notin\mathcal{M}_{s}$ for some $s\in C_{p^{\star}}$, then $\delta_{\nu}(s)>0$. Since $s$ lies in the recurrent class of $p^{\star}$, the time-average of $\delta_{\nu}(S_{t})$ under $(\nu,p^{\star})$ is strictly positive; hence by Lemma D.2, $g^{\nu}_{p^{\star}}<\alpha^{\star}$, which contradicts optimality (because $\min_{p}g^{\nu}_{p}\leq g^{\nu}_{p^{\star}}<\alpha^{\star}$).
(b) If $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for every $s$, then by (66) we have $u^{\star}(s)=\inf_{p_{s}}L_{s}\!\left(\nu(\cdot\mid s),p_{s}\right)\leq L_{s}\!\left(\nu(\cdot\mid s),p_{s}\right)$ for all $p_{s}$, so summation and telescoping give $g^{\nu}_{p}\geq\alpha^{\star}$ for every $p\in\mathcal{P}^{\pi_{-i}}$. On the other hand, $\alpha^{\star}$ is the optimal robust value for the induced DR-MDP (Bellman optimality under weak communication), so $\min_{p}g^{\nu}_{p}\leq\alpha^{\star}$. Therefore $\min_{p}g^{\nu}_{p}=\alpha^{\star}$, i.e., $\nu\in\text{BR}_{i}(\pi_{-i})$. ∎

### v3

@@ C.2 Hemi-continuity
Let $\{\pi_{n}\}$ be a sequence of stationary policies converging to $\pi$ in the product topology. Let $P^{\pi_{n}}$ and $P^{\pi}$ denote the induced kernels for a fixed $P\in\mathcal{P}$, and let $(P^{\pi_{n}})^{\star}$ and $(P^{\pi})^{\star}$ be their stationary distributions. Then
$(P^{\pi_{n}})^{\star}\;\to\;(P^{\pi})^{\star}$
uniformly in $P\in\mathcal{P}$.
@@
This is a direct consequence of the continuity properties of irreducible Markov chains; see (puterman2014markov) (Prop. 8.4.6). Uniformity in $P$ follows from the compactness of $\mathcal{P}$. ∎
@@
Fix an agent $i$ and let $\pi:\mathcal{S}\to\Delta(\mathcal{A}_{-i})$ be a joint policy of all other agents. Define the operator $T^{\pi}_{\mathcal{P}}:\mathbb{R}^{\mathcal{S}}\to\mathbb{R}^{\mathcal{S}\times\mathcal{A}_{i}}$ by
$T^{\pi}_{\mathcal{P}}h(s,a_{i})\;\triangleq\;r^{\pi}(s,a_{i})+\sigma^{\pi}_{s,a_{i}}(h),$
where
$r^{\pi}(s,a_{i})\triangleq\sum_{a_{-i}}\pi(a_{-i}\mid s)\,r_{i}\bigl(s,(a_{i},a_{-i})\bigr),$
and
$\sigma^{\pi}_{s,a_{i}}(h)\triangleq\min_{q\in\mathcal{P}^{\pi}_{s,a_{i}}}q^{\top}h,\qquad\mathcal{P}^{\pi}_{s,a_{i}}\triangleq\left\{\sum_{a_{-i}}\pi(a_{-i}\mid s)P(\cdot\mid s,(a_{i},a_{-i})):P(\cdot\mid s,(a_{i},a_{-i}))\in\mathcal{P}_{s}^{(a_{i},a_{-i})}\right\}.$
Let $\{\pi_{n}\}$ be a sequence of opponent policies with $\pi_{n}\to\pi$, and $\{h_{n}\}$ a sequence of functions $h_{n}\to h$ uniformly on $\mathcal{S}$. Then
$T^{\pi_{n}}_{\mathcal{P}}(h_{n})\;\to\;T^{\pi}_{\mathcal{P}}(h)$
uniformly on $\mathcal{S}\times\mathcal{A}_{i}$.
@@ Appendix D Weakly Communicating DR-MGs
Throughout this section we fix a player $i$ and a stationary joint policy $\pi_{-i}$ of the other players, and work with the induced single-agent DR-MDP
$M_{i}(\pi_{-i})\;=\;(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}}_{i}).$
All quantities defined below (transition sets, Bellman solutions, etc.) depend on $(i,\pi_{-i})$, but we often suppress this dependence to lighten notation.
Robust average reward and best responses.
For a stationary policy $\pi_{i}:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ of player $i$, denote by $g^{\pi_{i}}_{P}$ its average reward in $M_{i}(\pi_{-i})$ under a kernel $P\in\mathcal{P}^{\pi_{-i}}$. We define the robust worst-case average reward and the associated best-response set as
$g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}$ $\triangleq\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\pi_{i}}_{P},$ (48)
$\text{BR}_{i}(\pi_{-i})$ $\triangleq\Bigl\{\pi_{i}:g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}=\sup_{\mu:\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\mu}_{P}\Bigr\}.$ (49)
Constant-gain robust Bellman equations.
Under Assumption 2 the induced DR-MDP $M_{i}(\pi_{-i})$ is weakly communicating, so we may invoke the main result of (wang2025bellman).
Lemma D.1 (Constant-gain robust Bellman equations, strong duality).
For the induced DR-MDP $M_{i}(\pi_{-i})$ there exists a constant-gain solution $(u^{\star},g^{\star})$ to the state-wise min–max Bellman equations
$u^{\star}(s)$ $=\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\bigl(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+p_{s,a_{i}}^{\top}u^{\star}\bigr),$ (50)
$u^{\star}(s)$ $=\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\bigl(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+p_{s,a_{i}}^{\top}u^{\star}\bigr),\qquad s\in\mathcal{S},$ (51)
with attainment of both the $\inf$ and the $\sup$ for every state $s$. Moreover $g^{\star}$ equals the optimal robust average reward of $M_{i}(\pi_{-i})$,
$g^{\star}=\sup_{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\nu}_{P},$ (52)
and the sup in (52) is attained.
Lemma D.1 is a direct specialization of wang2025bellman to the induced DR-MDP $M_{i}(\pi_{-i})$. In particular, $(u^{\star},g^{\star})$ is a solution of the optimal robust Bellman equation for $M_{i}(\pi_{-i})$, and $g^{\star}$ is the optimal min–max value.
@@ D.1 Convexity
For each state $s\in\mathcal{S}$ define the local saddle Lagrangian
$L_{s}(\varphi,p_{s})~\triangleq~\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\bigl(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+p_{s,a_{i}}^{\top}u^{\star}\bigr),\quad\varphi\in\Delta(\mathcal{A}_{i}),\;p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}.$ (53)
We introduce the value functions
$F_{s}(\varphi)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}L_{s}(\varphi,p_{s}),$ (54)
$\Psi_{s}(p_{s})$ $\triangleq\sup_{\varphi\in\Delta(\mathcal{A}_{i})}L_{s}(\varphi,p_{s}).$ (55)
The function $F_{s}$ is concave and upper semicontinuous on the compact convex set $\Delta(\mathcal{A}_{i})$ (being the infimum of continuous affine maps), while $\Psi_{s}$ is convex and lower semicontinuous on $(\mathcal{P}^{\pi_{-i}})_{s}$.
By the duality relations (50)–(51),
@@
$M_{s}\;\triangleq\;\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F_{s}(\varphi),$ (57)
which is nonempty, convex, and compact by standard convex analysis (boyd2004convex).
@@
By Berge’s Maximum Theorem (berge1877topological), $\Psi_{s}$ is continuous on compact $(\mathcal{P}^{\pi_{-i}})_{s}$, so a minimizer $p_{s}^{\star}$ exists. Then it holds that $\Psi_{s}(p_{s}^{\star})=u^{\star}(s)$. For any $\varphi\in\mathcal{M}_{s}$,
@@
Slack inequalities and characterization of ${G}_{i}$.
For a stationary policy $\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ of player $i$ define the Bellman slack
$\delta_{\nu}(s)\;\triangleq\;u^{\star}(s)-F_{s}\bigl(\nu(\cdot\mid s)\bigr)\;\geq\;0,\qquad s\in\mathcal{S},$ (60)
where the inequality follows from (56).
Recall the definitions (for fixed $i$ and $\pi_{-i}$)
$L_{s}(\phi,p_{s})$ $\triangleq\sum_{a_{i}\in A_{i}}\phi(a_{i})\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Bigr),\quad\phi\in\Delta(A_{i}),\ p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s},$ (61)
$F_{s}(\phi)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}}L_{s}(\phi,p_{s}),\qquad\Psi_{s}(p_{s})\triangleq\sup_{\phi\in\Delta(A_{i})}L_{s}(\phi,p_{s}).$ (62)
By the duality relations (65)–(66), we have for every $s\in\mathcal{S}$
$u^{\star}(s)\;=\;\max_{\phi\in\Delta(A_{i})}F_{s}(\phi)\;=\;\min_{p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}}\Psi_{s}(p_{s}).$ (63)
Let
$M_{s}\;\triangleq\;\operatorname{argmax}_{\phi\in\Delta(A_{i})}F_{s}(\phi),$
which is nonempty, convex, and compact.
Lemma D.1 already shows that there exists a common supporting minimizer $p_{s}^{\star}\in(\mathcal{P}_{\pi_{-i}})_{s}$ such that
$u^{\star}(s)=\Psi_{s}(p_{s}^{\star})=L_{s}(\phi,p_{s}^{\star})=F_{s}(\phi),\quad\forall\,\phi\in M_{s}.$ (64)
By $(s,a_{i})$-rectangularity, $p^{\star}\triangleq\{p_{s}^{\star}\}_{s\in\mathcal{S}}\in\mathcal{P}_{\pi_{-i}}$. Under Assumption 2, this kernel admits at least one closed recurrent class $C_{p^{\star}}\subseteq\mathcal{S}$, and every state outside $C_{p^{\star}}$ is transient under any stationary policy. [^2] This is the standard decomposition of a finite weakly communicating Markov decision process into transient states and recurrent communicating classes; see, e.g., puterman2014markov.
@@
For a stationary policy $\nu:S\to\Delta(A_{i})$ define
$\delta_{\nu}(s)\;\triangleq\;u^{\star}(s)-F_{s}\bigl(\nu(\cdot\mid s)\bigr)\ \geq 0,\quad s\in\mathcal{S}.$
Let $(S_{t},A_{t})$ be the trajectory generated under $(\nu,p^{\star})$ starting from any initial distribution. Then
$\limsup_{n\to\infty}\;\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\bigr)\Bigr]$  $\;\leq\;0,$ (65)
$\liminf_{n\to\infty}\;\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\bigr)\Bigr]$ $\;\geq\;\liminf_{n\to\infty}\;\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr].$ (66)
Moreover, if $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$, then
$g^{\nu}_{p^{\star}}\;\triangleq\;\lim_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}r^{\pi_{-i}}_{i}(S_{t},A_{t})\Bigr]=\alpha^{\star}.$
@@
Step 1: upper bound. Fix any state $s\in\mathcal{S}$. By definition of $F_{s}$ and $u^{\star}(s)$ we have, for any action distribution $\phi\in\Delta(A_{i})$ and any $p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}$,
$L_{s}(\phi,p_{s})\;\geq\;F_{s}(\phi)\;\leq\;u^{\star}(s).$
In particular, plugging $\phi=\nu(\cdot\mid s)$ and the special kernel $p^{\star}_{s}$ from Lemma D.1 yields
$L_{s}\bigl(\nu(\cdot\mid s),p^{\star}_{s}\bigr)\;\leq\;\Psi_{s}(p^{\star}_{s})=u^{\star}(s).$
Writing this out,
$u^{\star}(s)\;\geq\;\sum_{a_{i}}\nu(a_{i}\mid s)\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p^{\star\top}_{s,a_{i}}u^{\star}\Bigr).$
Under $(\nu,p^{\star})$, $A_{t}\sim\nu(\cdot\mid S_{t})$ and the next state satisfies $\mathbb{E}[u^{\star}(S_{t+1})\mid S_{t}=s,A_{t}=a_{i}]=p^{\star\top}_{s,a_{i}}u^{\star}$. Therefore
$u^{\star}(S_{t})\;\geq\;\mathbb{E}\bigl[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\bigr]\quad\text{a.s.}$ (67)
Taking unconditional expectation and summing from $t=0$ to n1n-1 gives
$\mathbb{E}[u^{\star}(S_{0})]\;\geq\;\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\bigr)\Bigr].$
Rearrange the telescoping sum in $u^{\star}$:
$\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}r^{\pi_{-i}}_{i}(S_{t},A_{t})-n\alpha^{\star}\Bigr]\;\leq\;\mathbb{E}[u^{\star}(S_{0})]-\mathbb{E}[u^{\star}(S_{n})].$
Divide by $n$ and use boundedness of $u^{\star}$ to get
$\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\bigr)\Bigr]\;\leq\;\frac{\|u^{\star}\|_{\infty}+\|u^{\star}\|_{\infty}}{n}\to 0,$
which yields (73) after taking the $\limsup$.
Step 2: lower bound with slack. By definition of $\delta_{\nu}(s)$ and $F_{s}$,
$u^{\star}(s)-\delta_{\nu}(s)=F_{s}\bigl(\nu(\cdot\mid s)\bigr)=\inf_{p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}}\sum_{a_{i}}\nu(a_{i}\mid s)\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Bigr).$
In particular, evaluating the infimum at $p^{\star}_{s}$,
$u^{\star}(s)-\delta_{\nu}(s)\;\leq\;\sum_{a_{i}}\nu(a_{i}\mid s)\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p^{\star\top}_{s,a_{i}}u^{\star}\Bigr).$
As above, this translates to
$u^{\star}(S_{t})-\delta_{\nu}(S_{t})\;\leq\;\mathbb{E}\bigl[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\bigr]\quad\text{a.s.}$ (68)
Taking expectations and summing from $t=0$ to n1n-1,
$\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\bigl(u^{\star}(S_{t})-\delta_{\nu}(S_{t})\bigr)\Bigr]\;\leq\;\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\bigr)\Bigr].$
Rearranging,
$\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star})\Bigr]\;\geq\;\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr]+\mathbb{E}[u^{\star}(S_{0})]-\mathbb{E}[u^{\star}(S_{n})].$
Divide by $n$ and pass to the $\liminf$ to obtain (74).
Step 3: equality under greedy actions on $C_{p^{\star}}$. Assume now that $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$. Then, by (61)–(62) and Lemma D.1,
$F_{s}\bigl(\nu(\cdot\mid s)\bigr)=u^{\star}(s),\quad s\in C_{p^{\star}},$
hence $\delta_{\nu}(s)=0$ on $C_{p^{\star}}$.
Under $(\nu,p^{\star})$, states in $S\setminus C_{p^{\star}}$ are transient, and $C_{p^{\star}}$ is a closed recurrent class (by Assumption 2 and the standard structure of finite Markov chains). Thus the fraction of time spent in $S\setminus C_{p^{\star}}$ converges to zero almost surely, while any invariant distribution is supported on $C_{p^{\star}}$ only. It follows that
$\lim_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr]=0.$
Combining this with the upper and lower bounds (73)–(74) gives
$\lim_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}r^{\pi_{-i}}_{i}(S_{t},A_{t})\Bigr]=\alpha^{\star},$
i.e., $g^{\nu}_{p^{\star}}=\alpha^{\star}$. ∎
Proposition D.1 (Characterization of G{G}).
Define
${G}_{i}(\pi_{-i})\;\triangleq\;\Bigl\{\nu:\mathcal{S}\to\Delta(A_{i})\ :\ \nu(\cdot\mid s)\in M_{s}\ \text{for all }s\in C_{p^{\star}}\Bigr\},$
where $C_{p^{\star}}$ is any closed recurrent class of the kernel $p^{\star}$ constructed in Lemma D.1. Then:
  1. 1.
    If $\nu\in\mathrm{BR}_{i}(\pi_{-i})$, every closed recurrent class $C_{p^{\star}}$ of $p^{\star}$ satisfies $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$.
  2. 2.
    If $\nu\in\mathcal{G}_{i}(\pi_{-i})$, then $\nu\in\mathrm{BR}_{i}(\pi_{-i})$.
@@
(1) Suppose $\nu\in\mathrm{BR}_{i}(\pi_{-i})$ and let $C_{p^{\star}}$ be any closed recurrent class of $p^{\star}$. If there existed $s_{0}\in C_{p^{\star}}$ with $\nu(\cdot\mid s_{0})\notin M_{s_{0}}$, then $F_{s_{0}}(\nu(\cdot\mid s_{0}))<u^{\star}(s_{0})$, so $\delta_{\nu}(s_{0})>0$. Since $s_{0}$ lies in a recurrent class of $p^{\star}$, the chain $(S_{t})$ under $(\nu,p^{\star})$ visits $s_{0}$ with positive asymptotic frequency, implying
$\liminf_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr]>0.$
Lemma D.3 then yields
$\limsup_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star})\Bigr]<0,$
i.e., $g^{\nu}_{p^{\star}}<\alpha^{\star}$. On the other hand,
$\min_{p\in\mathcal{P}_{\pi_{-i}}}g^{\nu}_{p}\;\leq\;g^{\nu}_{p^{\star}}<\alpha^{\star},$
which contradicts the fact that $\alpha^{\star}$ is the optimal robust value $\sup_{\tilde{\nu}}\min_{p}g^{\tilde{\nu}}_{p}$ attained by $\nu\in\mathrm{BR}_{i}$. Hence no such $s_{0}$ can exist, and $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$.
(2) Conversely, assume $\nu\in\mathcal{G}_{i}(\pi_{-i})$, so $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$. By Lemma D.3, $g^{\nu}_{p^{\star}}=\alpha^{\star}$. For any $p\in\mathcal{P}_{\pi_{-i}}$, the same one-step inequality as in the proof of Lemma D.3 (but now with $p$ in place of $p^{\star}$ and again using $F_{s}(\nu(\cdot\mid s))\leq L_{s}(\nu(\cdot\mid s),p_{s})$) implies
$\liminf_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star})\Bigr]\;\geq\;0,$
so $g^{\nu}_{p}\geq\alpha^{\star}$ for all $p\in\mathcal{P}_{\pi_{-i}}$. By definition of $\alpha^{\star}$ as the optimal min–max value,
$\min_{p\in\mathcal{P}_{\pi_{-i}}}g^{\nu}_{p}\;\leq\;\sup_{\tilde{\nu}}\min_{p}g_{p}^{\tilde{\nu}}=\alpha^{\star},$
and therefore $\min_{p}g^{\nu}_{p}=\alpha^{\star}$, i.e., $\nu$ is a robust best response to $\pi_{-i}$. ∎